Let X be a continuous r.v with density function
f(x) = ( |x|
10
, for r 2 < x < 4 0, elsewhere
. Calculate the expected value of X.
1
Expert's answer
2020-05-11T14:47:14-0400

E(x)="\intop" x*f(x)dx="\intop" 42x*|x|/10=("\int" 42x2dx)/10=x3/30|42=28/15=1.86